# Going beyond Landauer scattering theory to describe spatially-resolved non-local heating and cooling in quantum thermoelectrics

### v1

@@ I Introduction
Dissipation is a natural by-product of electrical resistance, leading to Joule heating. This conversion of electric power into heat is often an inconvenience, but it is also exploited to produce heat and light (electric heaters and filament lamps). At the same time, some materials exhibit thermoelectric effects, that are exploited for applications such as refrigeration using Peltier cooling. These effects also exist at the nanoscale, where they are often understood using Landauer scattering theory that links charge transport [1, 2, 3, 4, 5, 6], thermal transport [2, 7] and thermoelectric effects [8, 9, 10, 11, 12], in a way that explains a variety of nanoscale experiments [13, 14, 15, 16]. However, there has long been a debate in the literature about where the dissipation actually occurs at the nanoscale. To this we can also add the question of where thermoelectric cooling occurs.
@@
The second phenomenon is directly related to where dissipation occurs, and is an unexpected consequence of the non-local nature of the heating and cooling of the phonons by the Joule heating and thermoelectric cooling induced by electron flow through the scatterer. If the heating and cooling of phonons were local, one would expect that heating and cooling would be maximal at the scatterer, and decaying away from the scatterer. Thus it is unexpected to find regimes where the maximal heating or cooling of the phonons is at a finite distance from the scatterer (as sketched in Fig. 1b), which we will refer to as heating and cooling spots. Indeed, in some cases, the effect even changes sign as one goes away from the barrier, with phonons being heated close to the scatterer but cooled further way (or vice-versa). We explain this unexpected phenomenon as follows; when an electron leaves the scatterer and enters the nanowire, that electron flies a certain distance before an electron-phonon scattering event, and this distance depends on that electron’s energy. Thus high-energy electrons and low-energy electrons thermalize with the phonons at different distances from the scatterer. This means that heating and cooling of phonons can occur at different distances from the scatterer.
While our model is minimal, and similar models have been treated elsewhere [28, 29, 30, 31], we believe the two phenomena that we identify will play a crucial role in local thermometry of nanowires containing scatterers such as a quantum dot. These could be InAs [14, 32, 33, 34] or other types of nanowires, as discussed in more detail in at the end of this paper. Of course, such experimental nanostructures have a host of parameters that our model does not include (the finite width of the nanowires, inelastic electron-electron scattering, and other scattering mechanisms), some of which are treated in other theoretical models [35, 36, 37, 38]. So we do not expect quantitative agreement between experiments and our minimal model. Nonetheless, the two phenomena that we identify appear robust enough that it is plausible that they will be present in many nanowires containing quantum dots. Thus, we expect they will be observable in scanning thermal microscopy of suitably chosen nanostructures.
@@ II.4 Heating and cooling of phonons
recalling that $\varphi(\pm\infty)=0$. Taking the sum of the two gives $P^{(\varphi)}_{\text{$\rm R$}}+P^{(\varphi)}_{\text{$\rm L$}}=I(\varphi(0^{+})-\varphi(0^{-}))$, which is the Joule heating for the voltage drop $(\varphi(0^{+})-\varphi(0^{-}))$. Thus the total Joule heating induced by the scatterer is
@@ IV.1 Why the left-right asymmetry strongly differs
The first difference between the full solution and the Landauer scattering theory prediction is physically straight-forward; the nanowires have a thermoelectric response, $J_{\text{el}}(\pm\infty)$ given by Eq. (15). This is absent in Landauer scattering theory, since such a Landauer scattering theory assumes leads with $\mu_{0}\gg k_{\text{B}}T$. Thus a nanowire containing a scatterer (as in Fig. 1) is actually three thermoelectrics in series. The scatterer is a thermoelectric that is sandwiched between two (semi-infinite) nanowires that are also thermoelectrics. The additive effect of the three thermoelectrics can be very different from that of the scatterer alone, particularly when the scatterer’s thermoelectric response is smaller than that of the two nanowires, so the thermoelectric response is dominated by the nanowires. With this in mind we make a simple improvement to Landauer scattering theory, we assume $J_{\text{el}}^{\rm Landauer}(0^{+})$ approximates the heat current leaving the scatterer at $x=0^{+}$, but then note that electrons deep in the wire (at $x\to\infty$) carry heat $J_{\text{el}}(\infty)$, so the heat flow into the phonons must be $J_{\text{el}}^{\rm Landauer}(0^{+})-J_{\text{el}}(\infty)$. Making the same argument for the left side of the nanowire, we would predict $P_{\mathrm{asym}}$ would be approximately
@@
Before discussing why this approximation is good in Fig. 7a but not Fig. 7b, we note that it allows us to understand why the Landauer scattering theory works better at low temperatures, and why the asymmetry changes sign as $T$ increases. The nanowires’ thermoelectric response (for given $I$) is always negligible at small $T$, and then grows with increasing $T$, as seen from Eq. (15). In contrast, the scatterer’s thermoelectric response (for given $I$) may be fairly weakly $T$ dependent. In that case, it is normal that small $T$ corresponds to a thermoelectric effect dominated by the scatterer, but large $T$ corresponds to one dominated by the nanowires. Hence if $P_{\rm asym}$ is positive at small $T$ (when it is dominated by the scatterer), one can guess that it will become negative whenever $T$ becomes large enough that the nanowire’s thermoelectric response starts to dominate.
@@
It then becomes easy to understand why the full theory — or the above approximation of it — can predict cooling of the phonons on the right of the scatterer, when Landauer predicts heating of those phonons (and vice versa for phonons on the left of the scatterer). A specific example of this is large $T$ in Fig. 7a. We already know that the nanowire’s Peltier effect induces a heat current in the same direction as the particle current, Eq. (15). For simplicity, let us assume the scatterer’s Peltier effect is of the same nature (typical of transmission that is peaked above $\mu_{0}$, as in Fig. 7). Thus when there is a particle current through all three thermoelectrics in series, their electrons all carry heat from left to right (cooling phonons to their left and heating phonons to their right). However, if the middle thermoelectric (the scatterer) exhibits a weaker Peltier effect than the nanowire to its right, then it will be heating phonons to its right by less than the right-hand thermoelectric (the right nanowire) is cooling those phonons. Thus the net effect is a cooling of the phonons to the right of the middle thermoelectric, the opposite of what one would think when considering the middle thermoelectric in isolation. Landauer scattering theory is effectively treating the middle thermoelectric in isolation, and forgetting the thermoelectric responses of the nanowires. This can only be justified when $T$ is low enough that the nanowires have a negligible thermoelectric responses. The approximation in Eq. (36) does much better by accounting for the nanowires’s thermoelectric responses.
@@
Now we turn to the second main difference between the full solution and the Landauer scattering theory prediction, that is not captured by the approximation in Eq. (36). In Landauer scattering theory, it is assumed that each lead is macroscopic, and so its electronic distribution is not perturbed by the scatterer or the current flowing through it (for example, the lead conductances are assumed to be infinitely greater than the scatterer’s conductance), which is why it takes the distribution in the leads to be $f_{0}$ from Eq. (7). Here, in contrast, the leads are nanowires with finite conductivities, so their electronic distributions are modified whenever they carry current. Each nanowire’s distribution of left and right movers is shifted in opposite directions, given by $\delta f_{\rm D}(p)$ in Eq. (12). Since $\delta f_{\rm D}(p)\propto l(p)\,\partial_{\mu_{0}}f_{0}(p)$, in addition to the thermal broadening, its peak shifts towards higher energies above the Fermi level by an amount $\sim{k}_{\text{B}}^{2}{T}^{2}/\mu_{0}$. The finite conductance of the nanowires also means there is a potential buildup, $\varphi(x)$, in the nanowire close to the scatterer, which not only gives the $P^{(\varphi)}_{\text{$\rm R$}}-P^{(\varphi)}_{\text{$\rm L$}}$ term in Eq. (34), it also further modifies the electronic distributions that enter $J_{\text{el}}(0^{\pm})$. Unfortunately, these factors are hard to guess and probably require the full self-consistent solution of the problem. However, in the weak transmission limit, $\mathcal{T}(\epsilon)\ll{1}$, these factors are of order $\mathcal{O}(\mathcal{T}^{2})$, since $\delta{f}_{D}(p)$ is $\mathcal{O}(\mathcal{T})$ with respect to $f_{0}(p)$, and $P^{(\varphi)}\sim\mathcal{O}(I\varphi(x))\sim\mathcal{O}(\mathcal{T}^{2})$. Thus the approximation $P_{\rm asym}^{\rm approx}$ captures all effects up to $\mathcal{O}(\mathcal{T})$, so it will become exact in the limit of small $\mathcal{T}$.
@@ IV.2 Non-locality of the phonon heating
Thus, to be rigorous, it is wise to separate heating at point $x$ into the part that is due to the voltage drop at $x$, and the part that is really non-local, because it is due to a voltage drop elsewhere. This separation is easily done. ${\cal P}_{\text{ph}}^{(\varphi)}(x)$ is local, because its value at $x$ depends on $\partial_{x}e\varphi(x)$. In contrast, ${\cal P}_{\text{ph}}^{\rm s}(x)$ is non-local because it is due to the voltage drop at the scatterer (at $x=0$). They have very different magnitudes for small currents, $I$. The non-local term ${\cal P}_{\text{ph}}^{\rm s}(x)$ is $\mathcal{O}[I]$, while the local term ${\cal P}_{\text{ph}}^{(\varphi)}(x)$ is $\mathcal{O}[I^{2}]$. Thus for small currents, the heating and cooling at $x\neq 0$ is entirely a consequence of the non-local part of the phonon heating. This is the case in Figs. 3 and 4.
@@ IV.3 Why the spatial profile has heating and cooling spots
The sketch in Fig. 8 is constructed for Eqs. (10), (26) as follows. The violet curve is the distribution of right-moving electrons leaving the scatterer at $x=0^{+}$. This distribution is made of a shifted Fermi function (representing reflection of left-moving electrons from the scatterer) with a peak super-imposed on it (representing right-moving electrons that were transmitted through the scatterer). As the right-moving electrons move away from the scatterer, the violet distribution relaxes towards the green distribution by exchanging energy with the phonons. This results in electrons at some energies heating phonons, and electrons at other energies cooling phonons (as indicated). However, when electrons have more energy, they move faster, and the heat exchange with phonons occurs over a longer distance. Thus heating and cooling effects happen on different length scales. In the example sketched here, the heating of phonons is dominated by slow (low energy) electrons, while cooling of phonons is dominated by fast (high energy) electrons. Hence, heating happens closer to the scatterer than cooling, as seen clearly to the right of the scatterer in Fig. 3a’s black curve.
@@ V Conditions for experimental observability of hot and cold spots
  • Inelastic electron-electron collisions in the nanowire are less frequent than electron-phonon collisions, so electrons do not thermalize between themselves before thermalizing with the phonons.
@@
The formulated conditions can be satisfied by samples such as InAs nanowires containing quantum dots, like those whose global thermoelectric response was studied at low temperatures in Ref. [14, 33], and at high temperatures in Refs. [32, 34, 45]. These nanowires’ electrochemical potential, $\mu_{0}$, can be tuned with doping or electro-static gates, and experimentally $\mu_{0}$ (as measured from the band-edge) is often of similar order to temperature, thus we take such parameters in our modeling. It means that the nanowire itself has a significant thermoelectric response (as observed experimentally in Ref. [14]).
To see if the heating and cooling spots could be resolved by scanning thermal microscopy (SThM), we should be aware that scanning thermal microscopy detects hot spots (local temperature increase of the phonons), not heating spots (local heat flow from electrons to phonons, as plotted in our figures). We can expect that the spatial variation of the phonon temperature will correspond to a broadened version of the spatial distribution of heating and cooling spots. This broadening will be of the order of the phonon mean free path. Thus, for the heating and cooling spots to be clearly visible in the phonon temperature profile, we need that the peaks of the spots are at a distance from the scatterer that is more than a phonon mean free path. Our results show that their distance from the scatterer is of order the electron-phonon scattering length. For semiconducting nanowires at 300 K this scattering length is typically hundreds of nanometers, as given in Table 1 of Ref. [46], while at low temperatures it is typically longer than at 300 K. In contrast, the phonon mean free path will typically be of order the wire diameter, which can be made 50 nm or smaller. In such cases, the heating and cooling spots will be visible despite the smoothing of the spatial profile of the phonon temperature.
@@ Acknowledgments
We gratefully acknowledge financial support from the French National Research Agency ANR through project ANR-20-CE30-0028 (TQT). We are extremely grateful to E. Zeldov for the stimulating discussions that encouraged us to carry out this research. We thank H. Linke for discussions on transport in nanowires, and thank G. Percebois and G. Weick for interesting and fruitful discussions.

### v2

@@ I Introduction
Dissipation is a natural by-product of electrical resistance, leading to Joule heating. This conversion of electric power into heat is often an inconvenience, but it is also exploited to produce heat and light (electric heaters and filament lamps). At the same time, some materials exhibit thermoelectric effects, that are exploited for applications such as refrigeration using Peltier cooling. These effects also exist at the nanoscale, where they are often understood using Landauer scattering theory that links charge transport [1, 2, 3, 4, 5, 6], thermal transport [2, 7] and thermoelectric effects [8, 9, 10, 11, 12], in a way that explains a variety of nanoscale experiments [13, 14, 15, 16, 17]. However, there has long been a debate in the literature about where the dissipation actually occurs at the nanoscale. To this we can also add the question of where thermoelectric cooling occurs.
@@
The second phenomenon is directly related to where dissipation occurs, and is an unexpected consequence of the non-local nature of the heating and cooling of the phonons by the Joule heating and thermoelectric cooling induced by electron flow through the scatterer. If the heating and cooling of phonons were more local, one would expect that heating and cooling would be maximal at the scatterer, and decay away from the scatterer. Indeed, this is seen for the hydrodynamic model in Ref. [29] (its Fig. 4 shows that while there may be complicated spatial variations within the scattering region, heating decays smoothly outside that region). Thus, it is unexpected to find regimes where the maximal heating or cooling of the phonons is at a finite distance from the scatterer (as sketched in Fig. 1b), which we will refer to as heating and cooling spots. Indeed, in some cases, the effect even changes sign as one goes away from the barrier, with phonons being heated close to the scatterer but cooled further way (or vice-versa). We explain this unexpected phenomenon as follows; when an electron leaves the scatterer and enters the nanowire, that electron flies a certain distance before an electron-phonon scattering event, and this distance depends on that electron’s energy. Thus, high-energy electrons and low-energy electrons thermalize with the phonons at different distances from the scatterer. This means that heating and cooling of phonons can occur at different distances from the scatterer.
While our model is minimal, and similar models have been treated elsewhere [30, 31, 32, 33], we believe the two phenomena that we identify will play a crucial role in local thermometry of nanowires containing scatterers such as a quantum dot. These could be InAs [14, 34, 35, 36, 17] or other types of nanowires, as discussed in more detail in at the end of this paper. Of course, such experimental nanostructures have a host of parameters that our model does not include (the finite width of the nanowires, inelastic electron-electron scattering, and other scattering mechanisms), some of which are treated in other theoretical models [37, 38, 39, 40, 29]. So we do not expect quantitative agreement between experiments and our minimal model. Nonetheless, the two phenomena that we identify appear robust enough that it is plausible that they will be present in many nanowires containing quantum dots. Thus, we expect they will be observable in scanning thermal microscopy of suitably chosen nanostructures.
@@ II.4 Heating and cooling of phonons
recalling that $\varphi(\pm\infty)=0$. Taking the sum of the two gives $P^{(\varphi)}_{\text{$\rm R$}}+P^{(\varphi)}_{\text{$\rm L$}}=I(\varphi(0^{+})-\varphi(0^{-}))$, which is the Joule heating for the voltage drop $(\varphi(0^{+})-\varphi(0^{-}))$. Thus, the total Joule heating induced by the scatterer is
@@ IV.1 Why the left-right asymmetry strongly differs
The first difference between the full solution and the Landauer scattering theory prediction is physically straight-forward; the nanowires have a thermoelectric response, $J_{\text{el}}(\pm\infty)$ given by Eq. (15). This is absent in Landauer scattering theory, since such a Landauer scattering theory assumes leads with $\mu_{0}\gg k_{\text{B}}T$. Thus, a nanowire containing a scatterer (as in Fig. 1) is actually three thermoelectrics in series. The scatterer is a thermoelectric that is sandwiched between two (semi-infinite) nanowires that are also thermoelectrics. The additive effect of the three thermoelectrics can be very different from that of the scatterer alone, particularly when the scatterer’s thermoelectric response is smaller than that of the two nanowires, so the thermoelectric response is dominated by the nanowires. With this in mind we make a simple improvement to Landauer scattering theory, we assume $J_{\text{el}}^{\rm Landauer}(0^{+})$ approximates the heat current leaving the scatterer at $x=0^{+}$, but then note that electrons deep in the wire (at $x\to\infty$) carry heat $J_{\text{el}}(\infty)$, so the heat flow into the phonons must be $J_{\text{el}}^{\rm Landauer}(0^{+})-J_{\text{el}}(\infty)$. Making the same argument for the left side of the nanowire, we would predict $P_{\mathrm{asym}}$ would be approximately
@@
Before discussing why this approximation is good in Fig. 7a but not Fig. 7b, we note that it allows us to understand why the Landauer scattering theory works better at low temperatures, and why the asymmetry changes sign as $T$ increases. The nanowires’ thermoelectric response (for given $I$) is always negligible at small $T$, and then grows with increasing $T$, as seen from Eq. (15). In contrast, the scatterer’s thermoelectric response (for given $I$) may be fairly weakly $T$ dependent. In that case, it is normal that small $T$ corresponds to a thermoelectric effect dominated by the scatterer, but large $T$ corresponds to one dominated by the nanowires. Hence, if $P_{\rm asym}$ is positive at small $T$ (when it is dominated by the scatterer), one can guess that it will become negative whenever $T$ becomes large enough that the nanowire’s thermoelectric response starts to dominate.
@@
It then becomes easy to understand why the full theory — or the above approximation of it — can predict cooling of the phonons on the right of the scatterer, when Landauer predicts heating of those phonons (and vice versa for phonons on the left of the scatterer). A specific example of this is large $T$ in Fig. 7a. We already know that the nanowire’s Peltier effect induces a heat current in the same direction as the particle current, Eq. (15). For simplicity, let us assume the scatterer’s Peltier effect is of the same nature (typical of transmission that is peaked above $\mu_{0}$, as in Fig. 7). Thus, when there is a particle current through all three thermoelectrics in series, their electrons all carry heat from left to right (cooling phonons to their left and heating phonons to their right). However, if the middle thermoelectric (the scatterer) exhibits a weaker Peltier effect than the nanowire to its right, then it will be heating phonons to its right by less than the right-hand thermoelectric (the right nanowire) is cooling those phonons. Thus, the net effect is a cooling of the phonons to the right of the middle thermoelectric, the opposite of what one would think when considering the middle thermoelectric in isolation. Landauer scattering theory is effectively treating the middle thermoelectric in isolation, and forgetting the thermoelectric responses of the nanowires. This can only be justified when $T$ is low enough that the nanowires have a negligible thermoelectric responses. The approximation in Eq. (36) does much better by accounting for the nanowires’s thermoelectric responses.
@@
Now we turn to the second main difference between the full solution and the Landauer scattering theory prediction, that is not captured by the approximation in Eq. (36). In Landauer scattering theory, it is assumed that each lead is macroscopic, and so its electronic distribution is not perturbed by the scatterer or the current flowing through it (for example, the lead conductances are assumed to be infinitely greater than the scatterer’s conductance), which is why it takes the distribution in the leads to be $f_{0}$ from Eq. (7). Here, in contrast, the leads are nanowires with finite conductivities, so their electronic distributions are modified whenever they carry current. Each nanowire’s distribution of left and right movers is shifted in opposite directions, given by $\delta f_{\rm D}(p)$ in Eq. (12). Since $\delta f_{\rm D}(p)\propto l(p)\,\partial_{\mu_{0}}f_{0}(p)$, in addition to the thermal broadening, its peak shifts towards higher energies above the Fermi level by an amount $\sim{k}_{\text{B}}^{2}{T}^{2}/\mu_{0}$. The finite conductance of the nanowires also means there is a potential buildup, $\varphi(x)$, in the nanowire close to the scatterer, which not only gives the $P^{(\varphi)}_{\text{$\rm R$}}-P^{(\varphi)}_{\text{$\rm L$}}$ term in Eq. (34), it also further modifies the electronic distributions that enter $J_{\text{el}}(0^{\pm})$. Unfortunately, these factors are hard to guess and probably require the full self-consistent solution of the problem. However, in the weak transmission limit, $\mathcal{T}(\epsilon)\ll{1}$, these factors are of order $\mathcal{O}(\mathcal{T}^{2})$, since $\delta{f}_{D}(p)$ is $\mathcal{O}(\mathcal{T})$ with respect to $f_{0}(p)$, and $P^{(\varphi)}\sim\mathcal{O}(I\varphi(x))\sim\mathcal{O}(\mathcal{T}^{2})$. Thus, the approximation $P_{\rm asym}^{\rm approx}$ captures all effects up to $\mathcal{O}(\mathcal{T})$, so it will become exact in the limit of small $\mathcal{T}$.
@@ IV.2 Non-locality of the phonon heating
Thus, to be rigorous, it is wise to separate heating at point $x$ into the part that is due to the voltage drop at $x$, and the part that is really non-local, because it is due to a voltage drop elsewhere. This separation is easily done. ${\cal P}_{\text{ph}}^{(\varphi)}(x)$ is local, because its value at $x$ depends on $\partial_{x}e\varphi(x)$. In contrast, ${\cal P}_{\text{ph}}^{\rm s}(x)$ is non-local because it is due to the voltage drop at the scatterer (at $x=0$). They have very different magnitudes for small currents, $I$. The non-local term ${\cal P}_{\text{ph}}^{\rm s}(x)$ is $\mathcal{O}[I]$, while the local term ${\cal P}_{\text{ph}}^{(\varphi)}(x)$ is $\mathcal{O}[I^{2}]$. Thus, for small currents, the heating and cooling at $x\neq 0$ is entirely a consequence of the non-local part of the phonon heating. This is the case in Figs. 3 and 4.
@@ IV.3 Why the spatial profile has heating and cooling spots
The sketch in Fig. 8 is constructed for Eqs. (10), (26) as follows. The violet curve is the distribution of right-moving electrons leaving the scatterer at $x=0^{+}$. This distribution is made of a shifted Fermi function (representing reflection of left-moving electrons from the scatterer) with a peak super-imposed on it (representing right-moving electrons that were transmitted through the scatterer). As the right-moving electrons move away from the scatterer, the violet distribution relaxes towards the green distribution by exchanging energy with the phonons. This results in electrons at some energies heating phonons, and electrons at other energies cooling phonons (as indicated). However, when electrons have more energy, they move faster, and the heat exchange with phonons occurs over a longer distance. Thus, heating and cooling effects happen on different length scales. In the example sketched here, the heating of phonons is dominated by slow (low energy) electrons, while cooling of phonons is dominated by fast (high energy) electrons. Hence, heating happens closer to the scatterer than cooling, as seen clearly to the right of the scatterer in Fig. 3a’s black curve.
@@ V Conditions for experimental observability of hot and cold spots
  • Electron energy relaxation is dominated by electron-phonon collisions, rather than inelastic electron-electron collisions (this is the opposite limit from Refs. [40, 29]).
@@
The formulated conditions could be satisfied by samples such as InAs nanowires containing quantum dots, like those whose global thermoelectric response was studied at low temperatures in Ref.[14, 35, 17], and at high temperatures in Refs. [34, 36, 47]. Similar materials appear to have electron energy relaxation dominated by electron-phonon interactions above about 10 K [48]. These nanowires’ electrochemical potential, $\mu_{0}$, can be tuned with doping or electro-static gates, and experimentally $\mu_{0}$ (as measured from the band-edge) is often of similar order to temperature, thus we take such parameters in our modeling. It means that the nanowire itself has a significant thermoelectric response (as observed experimentally in Ref. [14]).
To see if the heating and cooling spots could be resolved by scanning thermal microscopy (SThM), we should be aware that scanning thermal microscopy detects hot spots (local temperature increase of the phonons), not heating spots (local heat flow from electrons to phonons, as plotted in our figures). We can expect that the spatial variation of the phonon temperature will correspond to a broadened version of the spatial distribution of heating and cooling spots. This broadening will be of the order of the phonon mean free path. Thus, for the heating and cooling spots to be clearly visible in the phonon temperature profile, we need that the peaks of the spots are at a distance from the scatterer that is more than a phonon mean free path. Our results show that their distance from the scatterer is of order the electron-phonon scattering length. For semiconducting nanowires at 300 K this scattering length is typically hundreds of nanometers [49], while at low temperatures it is typically longer than at 300 K. In contrast, the phonon mean free path will typically be of order the wire diameter, which can be made 50 nm or smaller. In such cases, the heating and cooling spots will be visible despite the smoothing of the spatial profile of the phonon temperature.
@@ Acknowledgments
We gratefully acknowledge financial support from the French National Research Agency ANR through project ANR-20-CE30-0028 (TQT). This research was funded by the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and DIPC on behalf of the Department of Education of the Basque Government. We are extremely grateful to E. Zeldov for the stimulating discussions that encouraged us to carry out this research. We thank H. Linke for discussions on transport in nanowires, and thank G. Percebois and G. Weick for interesting and fruitful discussions.